OpenStudy (anonymous):

can u plz help me solve the inequality -|1-2 x |≥-3

OpenStudy (anonymous):

-(1+2x) is what you're solving basically.

OpenStudy (anonymous):

*solving for, solve for x after distributing in the negative

OpenStudy (anonymous):

-\[-\left| 1-2x \right|\ge-3,\left| 1-2x \right|\le3,-3\le \left( 1-2x \right)\le3\] -4<=(1-2x-1)<=3-1 -4<=-2x<=2 4>=2x>=-2 2>=x>=-1 or -1<=x<=2 or x is in [-1,2],[ ] =closed interval

OpenStudy (anonymous):

thank u!!!